OpenStudy (anonymous):

93 /100 converted to percent equals?

OpenStudy (anonymous):

93%

myininaya (myininaya):

\[\frac{a}{100}=a \%\]

OpenStudy (karatechopper):

93%

OpenStudy (karatechopper):

whenever givin a fraction over 100 to change to percent, move the numerator's decimal 2 steps to the left, and tht will give u your percent!!

myininaya (myininaya):

\[\frac{\frac{a}{b}}{100}=\frac{a}{b} \%\]